# Dyons in higher-dimensional gauge theories

### v1

@@ Abstract
We discuss the ’t Hooft-Polyakov (TP) monopole and then dyon in the framework of the higher dimensional gauge theories, such as the gauge-Higgs unification models. First, we point out that the Bogomol’nyi-Prasad-Sommerfield (BPS) monopole is nothing but a self-dual gauge field in the 4-dimensional (4D) space including the extra dimension. This property is argued to have an inevitable remarkable consequence that the mass of the BPS monopole $M_{{\rm TPM}}$ and therefore the vacuum expectation value (VEV) of the Higgs field are topologically quantized. Next, the argument is generalized to the case of dyon.The mass of the BPS dyon, $M_{{\rm BPS}}$, is still proportional to the quantized Higgs VEV, though it also depends on a parameter $\mu$, denoting the ratio of the electric and magnetic charges of the dyon. In this paper, we propose a novel mechanism to quantize the parameter $\mu$ too. We demonstrate that in the 5D gauge theories the Chern-Simons term is induced at the quantum level, which, after the extra space component of the gauge field is replaced by its VEV, produces the $\theta$ term. Here again the VEV is quantized and through the Witten effect we reach to a remarkable conclusion that the electric charge of the dyon, and therefore the parameter $\mu$ is discretized. This means that $M_{{\rm BPS}}$ is also discretized. In addition, we propose a numerical method to obtain the field configurations and the mass of the non-BPS dyons by use of “modified” gradient flow equations. We find that the obtained field configurations of non-BPS dyon and its mass are rather close to those of the BPS dyon.
@@ 1 Introduction
First, we discuss that, as a specific feature of higher dimensional gauge theory, the Bogomol’nyi-Prasad-Sommerfield (BPS) saturated TP monopole (say, BPS monopole), which satisfies the BPS condition and therefore the equation of motion justified for the case of negligible Higgs potential, is nothing but an (anti-)self-dual gauge field in the 4D space with coordinates $(\vec{x},y)$ including the extra space (not 4D space-time) [18]. This is not just the matter of interpretation, but it is demonstrated to have a remarkable consequence. Namely, just as the action of the instanton, an (anti-)self-dual gauge field in the ordinary 4D space-time, is quantized ($\pi_{3}(SU(2))=Z$), the mass of the BPS monopole $M_{{\rm TPM}}$ (TPM stands for the TP monopole) is known to be quantized, being described just by the higher dimensional gauge coupling $g$ as $M_{{\rm TPM}}=\frac{8\pi^{2}}{g^{2}}$, corresponding to the Pontryagin index $\nu=-1$ (see eq.(11)). Or, in other words, the VEV of the Higgs field, or equivalently that of $A_{y}$, is quantized to a specific value from the 4D point of view. This is in clear contrast to the case of ordinary 4D space-time where $M_{{\rm TPM}}$ is proportional to the Higgs VEV, which is left as a free parameter, since the Higgs potential to fix the VEV is absent for the BPS monopole.
@@
In this paper, we propose a novel mechanism to fix $\mu$, which again utilizes characteristic features of higher dimensional gauge theories. Let us note that dyons, possessing both of electric and magnetic charges, may be understood to be P and CP-violating field configurations. Witten has proposed an interesting way to fix the electric charge $q$ of dyons in the theory containing the $\theta$ term, $\theta F_{\mu\nu}\tilde{F}^{\mu\nu}$, which clearly violates both of P and CP symmetries: it has been shown that $q$ is fixed in proportion to $\theta$ [20]. In [20], however, the origin of the $\theta$ term was not specified. We propose an approach to attribute the origin to the radiatively induced Chern-Simons (CS) term of the form $C_{y}F_{\mu\nu}\tilde{F}^{\mu\nu}$ in a 5D gauge theory, where $C_{y}$ denotes the extra space component of additional U(1) gauge field, independent of $A_{\mu},\ A_{y}$ to form the dyon solution. Then the VEV of $C_{y}$ clearly behaves as $\theta$. A remarkable feature of this approach is that the VEV of $C_{y}$ is quantized, just as in the case of the VEV of $A_{y}$. As the result, $\theta$ and therefore $q$ and finally $\mu$ are quantized through the “Witten effect”, a feature not shared by ordinary 4D theories. Thus we reach to an interesting conclusion that the mass of the dyon only takes discrete values.
@@
In this paper, we adopt the method of gradient flow to obtain the field configuration and the mass of the dyon numerically. For instance, in the simplified case of the BPS monopole (i.e. $\mu=0$), it is shown that by utilizing the gradient of the Hamiltonian in the gradient flow equation, we can readily reconstruct the analytic solutions correctly. In the case of the dyon ($\mu\neq 0$), however, because of the contribution of the “electric” field (though the gauge group is non-Abelian), the extremum conditions of lagrangian $L$, or rather $-L$, and the Hamiltonian $H$ do not coincide, and the gradient of $-L$ (not $H$) should be adopted to get the solution of the equation of motion correctly. However, such approach turns out to face a problem: it causes an unstable behavior in the flow of the field configuration. The reason for the instability should be the fact that the contribution of the “electric” field to $-L$ is semi-negative-definite.
@@ 2 The BPS monopole as a self-dual field in 5D gauge theory
Before discussing dyons, we first argue that the BPS saturated TP monopole, say BPS monopole, is able to be understood as an (anti-)self-dual gauge field in 5D SU(2) GHU [18]. It should be noted that in the 5D theory, the adjoint Higgs field $\phi$, indispensable for the construction of TP monopole has been built in as the extra space component of the 5D gauge field, $A_{y}$.
@@
is the well-known quantized magnetic charge and $g_{4}$, $v_{4}$ are the 4D gauge coupling and the magnitude of the VEV of $\phi$, related to the 5D counterparts $g,\ v$ by multiplying suitable powers of the factor $\sqrt{2\pi R}$. In the ordinary 4D space-time, there is no way to fix the mass $M_{{\rm TPM}}$, since $v_{4}$ is left as a free parameter. Note that the potential of $\phi$, to fix the VEV, is assumed to be absent for the case of BPS monopole.
In the framework of 5D GHU, however, since the BPS monopole is interpreted as a self-dual gauge field, we naturaly expect that its mass cannot be arbitrary but is topologically quantized, just as the action of the instanton is quantized in 4D space-time.
@@ 3 Dyon as a self-dual-like field in 5D gauge theory
$-2\int\ {\rm Tr}\left(B_{i}F_{iy}\right)\ {\rm d}^{3}x{\rm d}y=-2\int\ {\rm Tr}\left(B_{i}D_{i}A_{y}\right)\ {\rm d}^{3}x{\rm d}y=-2\int\ {\rm Tr}\partial_{i}\left(B_{i}A_{y}\right){\rm d}^{3}x{\rm d}y=v_{4}g_{m}.$ (23)
@@
Now the remaining constraint (17), i.e. $F_{0y}=0$, is automatically satisfied since fields are static KK zero modes and also $[A_{0},A_{y}]=0$ under (37). Thus we have obtained the field configuration of BPS dyon, utilizing the BPS monopole solution, modified by the angle $\mu$, which is so far left as an arbitrary parameter.
@@
One remark here is that, as we have already used the fact, since the magnetic charge $g_{m}$ is topologically quantized, we should have the same quantized value $g_{m}=\frac{4\pi}{g_{4}}$ as in the case of TP monopole, although the argument $\rho$ of $F(\rho),\ G(\rho)$ has been rescaled by the factor $\cos\mu$. This can be explicitly shown by calculating a surface integral of the magnetic flux $-\frac{1}{v}2{\rm Tr}(A_{y}B_{i})$, written in terms of $F,G$, on the sphere ${\rm S}^{2}$ located at spatial infinity $r\to\infty$:
@@ 4 Induced Chern-Simons term and quantized electric charge of dyon
An important novel result is that the dyon mass is quantized as well. Namely, by use of (11) and (30) the mass of the BPS dyon is expressed just by the radius of the compactified circle $R$ and the 4D gauge coupling $g_{4}$,
@@ 5 The analysis of dyon solution based on gradient flow equations
where $H$ is the Hamiltonian of the theory and $\Phi$ stands for a generic function to describe gauge field configuration, which is supposed to be not only $\rho$ dependent but also depends on the newly introduced variable $s$: $\Phi=\Phi(s,\rho)$. The essential point in this method is that the energy of the system decreases as $s$ increases and for $s\ \to\ \infty$, the fields approaches to the configurations satisfying $\frac{\delta H}{\delta\Phi}=0$, which realize the minimum energy. In this method, the boundary conditions at both ends, $\rho=0$ and $\rho=\infty$, are always kept during the $s-$evolution.
For instance, in the case of BPS monopole, corresponding to $\mu=0$ and therefore $J=0$ (see (55)), $\Phi=F,\ G$ and the Hamiltonian is given in terms of $F,\ G$ as,
@@
where $F^{\prime}\equiv\frac{{\rm d}F(\rho)}{{\rm d}\rho},\ G^{\prime}\equiv\frac{{\rm d}G(\rho)}{{\rm d}\rho}$. Let us note that in this case, because of the subsidiary constraint (6), i.e. $E_{i}=0$, the lagrangian $L$ is simply related to $H$ as $H=-L$, and the minimization of $H$, $\frac{\delta H}{\delta\Phi}=0$ just provides the solution to the equations of motion, $\frac{\delta L}{\delta\Phi}=0$.
In fact, as is seen in Fig.1, we have confirmed that the solutions for the gradient flow equations with respect to $F,\ G$ correctly reproduce the functions shown in (55) and also the correct value of the monopole mass in the limit of sufficiently large $s$, with the boundary conditions, $F(0)=G(0)=0$ and $F(\infty)=G(\infty)=1$, being imposed.
@@
In the case of the dyon, however, the situation is a little complicated, since in this case $E_{i}$ should exist inevitably and $H=-L$ no longer holds, even though the fields are all static. Namely, ${\cal H}={\rm Tr}\ \left[E_{i}^{2}+B_{i}^{2}+F_{iy}^{2}+V(A_{y})\right]$ and ${\cal L}={\rm Tr}\ \left[E_{i}^{2}-B_{i}^{2}-F_{iy}^{2}-V(A_{y})\right]$ leads to a relation
@@
On the other hand, it may be worth noting that, under the Gauss law constraint $D_{i}E_{i}=0$ (Eq.(18)), which is nothing but one of the equations of motion, the difference between $H$ and $-L$ turns out to be a surface term, and becomes a constant once the boundary conditions of the fields are fixed. To show that, first let us note that, since $E_{i}=-D_{i}A_{0}$,
@@
In fact, utilizing the constraint, which is explicitly written in terms of $J,\ G$ as
@@
We also get $J(\infty)=\sin\mu$ as in the case of BPS dyon, since for $\rho\to\infty$ the potential vanishes and the asymptotic behaviour of the fields should be the same as in the case of the BPS dyon,
@@
Hence, if we impose the constraint (60), relating $J$ with $G$, and regard $H$ and $-L$ as functionals of two independent functions, e.g. $G$ and $F$, the minimizations of $H$ and $-L$ should result in the same field configurations, since the surface term does not contribute to the functional derivatives.
However, when $H$ and $-L$ are regarded as the functionals of three independent functions, $F,\ G$ and $J$, we should be careful, since the extremum conditions $\frac{\delta H}{\delta\Phi}=0$ and $\frac{\delta(-L)}{\delta\Phi}=0$ for $\Phi=F,G,J$ yield different results. To be more concrete, $\frac{\delta H}{\delta\Phi}=0$ does not yield the solutions to the equations of motion, as we have already pointed out.
Since we are interested in the solutions to the equations of motion, our task seems to solve the gradient flow equations with respect to the three independent functions, $F,G$ and $J$, where $-L$ takes place of $H$. By the way, only in solving the gradient flow equation, by some technical reason in the numerical computation, we adopt $H(\rho),I(\rho)$ instead of $F(\rho),J(\rho)$, defined by
@@
Unfortunately, by solving the gradient flow equations (68) to (70) numerically , we have found that the function $I$ shows divergent behavior as $s$ increases. The origin of this problem seems to come from the fact that in the $-L$ the term $-E_{i}^{2}$ has a negative metric, which, when $-L$ is written in the quadratic form of the small deviations $\delta\Phi\ (\Phi=I,G,H)$ from the stationary point $\Phi_{st}$ (for which the r.h.s.s of (68) to (70) vanish) as $\Phi=\Phi_{st}+\delta\Phi$, probably leads to a negative eigenvalue of the matrix to describe the quadratic form and therefore the instability of the stationary point.
In order to overcome this problem, since the problematic term $-E_{i}^{2}$ is mainly described by $I$, we think about the possibility to evade the equation (68) and replace it by imposing the analytic Gauss law constraint (60) , now written in terms of $I$ as
@@
for the purpose to express $I$ in terms of $G$ and substitute such obtained $I$ in the r.h.s. of the remaining gradient flow equation (69). Nice thing of this method is once $I$ is enforced to be related to $G$ by use of the constraint, $-L$ should not have the problem of negative eigenvalue mentioned above. This is because, under the Gauss law constraint, $-L$ differs from $H$ just by a constant (as is seen in (66)), while $H$ is clearly positive definite and does not possess the problem of negative eigenvalue.
@@
+$+\left\{H^{\prime}-\sin\mu I^{\prime}-\frac{H-\sin\mu I}{\rho}+\cos^{2}\mu\frac{G(G-2)}{\rho}\right\}^{2}+\frac{2\rho^{2}}{g_{4}^{2}v_{4}^{4}\cos^{2}\mu}V_{4}(H/\rho)\Bigg{]},$

### v2

@@ Abstract
We discuss the ’t Hooft-Polyakov (TP) monopole and then dyon in the framework of higher dimensional gauge theories, such as gauge-Higgs unification models. First, we point out that the Bogomol’nyi-Prasad-Sommerfield (BPS) monopole is nothing but a self-dual gauge field in the 4-dimensional (4D) space including the extra dimension, which is argued to lead to a consequence that the mass of the BPS monopole $M_{{\rm TPM}}$ and therefore the vacuum expectation value (VEV) of the Higgs field are topologically quantized. In literatures, there exist related arguments on the calorons, which may be understood to be a composition of a pair of constituent monopole and anti-monopole, with each constituent carrying fractional topological charge, while the net topological charge carried by the caloron is unity. From the viewpoint of the caloron, our conclusion of the quantized monopole mass corresponds to the special case, where only a single monopole exists that carries the net topological charge. Next, the argument is generalized to the case of dyon. The mass of the BPS dyon, $M_{{\rm BPS}}$, is still proportional to the quantized Higgs VEV, though it also depends on a parameter $\mu$, denoting the ratio of the electric and magnetic charges of the dyon. In the 5D gauge theories the Chern-Simons term is induced at the quantum level, which, after the extra space component of the gauge field is replaced by its VEV, produces the $\theta$ term. Then, through the Witten effect we reach to an interesting conclusion that the parameter $\mu$ and therefore $M_{{\rm BPS}}$ are discretized. In addition, we propose a numerical method to obtain the field configurations and the mass of the non-BPS dyons by use of “modified” gradient flow equations.
@@ 1 Introduction
First, we discuss that, as a specific feature of higher dimensional gauge theory, the Bogomol’nyi-Prasad-Sommerfield (BPS) saturated TP monopole (say, BPS monopole), which satisfies the BPS condition and therefore the equation of motion justified for the case of negligible Higgs potential, is nothing but an (anti-)self-dual gauge field in the 4D space with coordinates $(\vec{x},y)$ including the extra space (not 4D space-time) [18]. This is not just the matter of interpretation, but it is demonstrated to have an interesting consequence. Namely, just as the action of the instanton, an (anti-)self-dual gauge field in the ordinary 4D space-time, is quantized ($\pi_{3}(SU(2))=Z$), the mass of the BPS monopole $M_{{\rm TPM}}$ (TPM stands for the TP monopole) is known to be quantized, being described just by the higher dimensional gauge coupling $g$ as $M_{{\rm TPM}}=\frac{8\pi^{2}}{g^{2}}$, corresponding to the Pontryagin index $\nu=-1$ (see eq.(11)). Or, in other words, the VEV of the Higgs field, or equivalently that of $A_{y}$, is quantized to a specific value from the 4D point of view. This is in clear contrast to the case of ordinary 4D space-time where $M_{{\rm TPM}}$ is proportional to the Higgs VEV, which is left as a free parameter, since the Higgs potential to fix the VEV is absent for the BPS monopole.
In literatures, there are related discussions concerning calorons, which are instanton-like self-dual gauge fields, that are periodic in the compactified Euclidean time [19, 20, 21]. Interestingly, it has been shown that a caloron can be understood to be a composition of a pair of constituent monopole and anti-monopole , where each constituent carries fractional topological charge, while the net topological charge carried by the caloron itself is unity [20]. As is discussed in more details in the next section, from the viewpoint of the caloron, our conclusion of the quantized monopole mass corresponds to the specific case, where only a single monopole exists that carries the net integer topological charge.
@@
In this paper, we propose a mechanism to fix $\mu$, which again utilizes characteristic features of higher dimensional gauge theories. Let us note that dyons, possessing both of electric and magnetic charges, may be understood to be P and CP-violating field configurations. Witten has proposed an interesting way to fix the electric charge $q$ of dyons in the theory containing the $\theta$ term, $\theta F_{\mu\nu}\tilde{F}^{\mu\nu}$, which clearly violates both of P and CP symmetries: it has been shown that $q$ is fixed in proportion to $\theta$ [22]. In [22], however, the origin of the $\theta$ term was not specified. We propose an approach to attribute the origin to the radiatively induced Chern-Simons (CS) term of the form $C_{y}F_{\mu\nu}\tilde{F}^{\mu\nu}$ in a 5D gauge theory, where $C_{y}$ denotes the extra space component of additional U(1) gauge field, independent of $A_{\mu},\ A_{y}$ to form the dyon solution. Then the VEV of $C_{y}$ clearly behaves as $\theta$. A characteristic feature of this approach is that the VEV of $C_{y}$ is quantized, just as in the case of the VEV of $A_{y}$. As the result, $\theta$ and therefore $q$ and finally $\mu$ are quantized through the “Witten effect”, a feature not shared by ordinary 4D theories. Thus we reach to an interesting conclusion that the mass of the dyon only takes discrete values.
@@
In this paper, we adopt the method of gradient flow to obtain the field configuration and the mass of the dyon numerically. For instance, in the simplified case of the BPS monopole (i.e. $\mu=0$), it is shown that by utilizing the gradient of the Hamiltonian in the gradient flow equation, we can readily reconstruct the analytic solutions correctly. In the case of the dyon ($\mu\neq 0$), however, because of the contribution of the “electric” field (though the gauge group is non-Abelian), the extremum conditions of lagrangian $L$, or rather L-L, and the Hamiltonian $H$ do not coincide, and the gradient of L-L (not $H$) should be adopted to get the solution of the equation of motion correctly. However, such approach turns out to face a problem: it causes an unstable behavior in the flow of the field configuration. The reason for the instability should be the fact that the contribution of the “electric” field to L-L is semi-negative-definite.
@@ 2 The BPS monopole as a self-dual field in 5D gauge theory
Before discussing dyons, we first argue that the BPS saturated TP monopole, say BPS monopole, is able to be understood as an (anti-)self-dual gauge field in 5D SU(2) GHU [18]. [^2] Recently, cosmic string, which is another solution of the soliton, has been studied in a similar setup [23]. It should be noted that in the 5D theory, the adjoint Higgs field $\phi$, indispensable for the construction of TP monopole has been built in as the extra space component of the 5D gauge field, $A_{y}$.
@@
is the well-known quantized magnetic charge and $g_{4}$, $v_{4}$ are the 4D gauge coupling and the magnitude of the VEV of $\phi$, related to the 5D counterparts g,vg,\ v by multiplying suitable powers of the factor $\sqrt{2\pi R}$. In the ordinary 4D space-time, there is no way to fix the mass $M_{{\rm TPM}}$, since $v_{4}$ is left as a free parameter. Note that the potential of $\phi$, to fix the VEV, is assumed to be absent for the case of BPS monopole.
In the framework of 5D GHU, however, since the BPS monopole is interpreted as a self-dual gauge field, we naturally expect that its mass cannot be arbitrary but is topologically quantized, just as the action of the instanton is quantized in 4D space-time.
@@
In literatures, there exist arguments on the related topics, i.e. calorons, which are instanton-like (anti-)self-dual gauge fields, that are periodic in the compactified Euclidean time, with a period $\beta=\frac{1}{T}$ denoting the inverse temperature in the context of finite temperature gauge theory [19, 20, 21]. Though calorons are periodic instantons rather than monopoles, interestingly, in ref.[20] it has been shown that a caloron may be understood to be a composition of a pair of constituent monopole and anti-monopole with opposite magnetic charges. In the argument, each constituent is regarded to carry fractional topological charge $k_{1}$ or $k_{2}$ (corresponding to the Pontryagin index) handled by $u$, the VEV of $A_{4}$ (corresponding to $gA_{y}$ in our case),
$k_{1}=\frac{\beta u}{2\pi},\ \ k_{2}=1-\frac{\beta u}{2\pi},$ (12)
though the net topological charge carried by the caloron is unity, $k_{1}+k_{2}=1$.
The difference of our argument from the argument for the caloron given above is that in our case the self-dual field is for a single monopole, not for a pair of monopoles, which itself carries the integer topological charge $|\nu|=1$. From the viewpoint of the caloron, the situation in our paper corresponds to the specific case $u=\frac{2\pi}{\beta}$, i.e. $k_{1}=1,\ k_{2}=0$. As is discussed below, in the GHU scenario, the VEV of $A_{y}$ cannot be arbitrary but is shown to be discretized relying on a general argument.
@@ 3 Dyon as a self-dual-like field in 5D gauge theory
$2\int\ {\rm Tr}\left(B_{i}F_{iy}\right)\ {\rm d}^{3}x{\rm d}y=-2\int\ {\rm Tr}\left(B_{i}D_{i}A_{y}\right)\ {\rm d}^{3}x{\rm d}y=-2\int\ {\rm Tr}\partial_{i}\left(B_{i}A_{y}\right){\rm d}^{3}x{\rm d}y=v_{4}g_{m}.$ (24)
@@
Now the remaining constraint (18), i.e. $F_{0y}=0$, is automatically satisfied since fields are static KK zero modes and also $A_{0},A_{y}]=0$ under (38). Thus we have obtained the field configuration of BPS dyon, utilizing the BPS monopole solution, modified by the angle $\mu$, which is so far left as an arbitrary parameter.
@@
One remark here is that, as we have already used the fact, since the magnetic charge $g_{m}$ is topologically quantized, we should have the same quantized value $g_{m}=\frac{4\pi}{g_{4}}$ as in the case of TP monopole, although the argument $\rho$ of $F(\rho),\ G(\rho)$ has been rescaled by the factor $\cos\mu$. This can be explicitly shown by calculating a surface integral of the magnetic flux $\frac{1}{v}2{\rm Tr}(A_{y}B_{i})$, written in terms of $F,G$, on the sphere ${\rm S}^{2}$ located at spatial infinity $r\to\infty$:
@@ 4 Induced Chern-Simons term and quantized electric charge of dyon
An important result is that the dyon mass is quantized as well. Namely, by use of (11) and (31) the mass of the BPS dyon is expressed just by the radius of the compactified circle $R$ and the 4D gauge coupling $g_{4}$,
@@ 5 The analysis of dyon solution based on gradient flow equations
where $H$ is the Hamiltonian of the theory and $\Phi$ stands for a generic function to describe gauge field configuration, which is supposed to be not only $\rho$ dependent but also depends on the newly introduced variable $s$: $\Phi=\Phi(s,\rho)$. The essential point in this method is that the energy of the system decreases as $s$ increases and for ss\ $\to\ \infty$, the fields approaches to the configurations satisfying $\frac{\delta H}{\delta\Phi}=0$, which realize the minimum energy. In this method, the boundary conditions at both ends, $\rho=0$ and $\rho=\infty$, are always kept during the $s$-evolution.
For instance, in the case of BPS monopole, corresponding to $\mu=0$ and therefore $J=0$ (see (56)), $\Phi=F,\ G$ and the Hamiltonian is given in terms of F,GF,\ G as,
@@
where $F^{\prime}\equiv\frac{{\rm d}F(\rho)}{{\rm d}\rho},\ G^{\prime}\equiv\frac{{\rm d}G(\rho)}{{\rm d}\rho}$. Let us note that in this case, because of the subsidiary constraint (6), i.e. $E_{i}=0$, the lagrangian $L$ is simply related to $H$ as H=LH=-L, and the minimization of $H$, $\frac{\delta H}{\delta\Phi}=0$ just provides the solution to the equations of motion, $\frac{\delta L}{\delta\Phi}=0$.
In fact, as is seen in Fig.1, we have confirmed that the solutions for the gradient flow equations with respect to F,GF,\ G correctly reproduce the functions shown in (56) and also the correct value of the monopole mass in the limit of sufficiently large $s$, with the boundary conditions, $F(0)=G(0)=0$ and $F(\infty)=G(\infty)=1$, being imposed.
@@
In the case of the dyon, however, the situation is a little complicated, since in this case $E_{i}$ should exist inevitably and H=LH=-L no longer holds, even though the fields are all static. Namely, ${\cal H}={\rm Tr}\ \left[E_{i}^{2}+B_{i}^{2}+F_{iy}^{2}+V(A_{y})\right]$ and ${\cal L}={\rm Tr}\ \left[E_{i}^{2}-B_{i}^{2}-F_{iy}^{2}-V(A_{y})\right]$ leads to a relation
@@
On the other hand, it may be worth noting that, under the Gauss law constraint $D_{i}E_{i}=0$ (Eq.(19)), which is nothing but one of the equations of motion, the difference between $H$ and L-L turns out to be a surface term, and becomes a constant once the boundary conditions of the fields are fixed. To show that, first let us note that, since $E_{i}=-D_{i}A_{0}$,
@@
In fact, utilizing the constraint, which is explicitly written in terms of J,GJ,\ G as
@@
We also get $J(\infty)=\sin\mu$ as in the case of BPS dyon, since for $\rho\to\infty$ the potential vanishes and the asymptotic behavior of the fields should be the same as in the case of the BPS dyon,
@@
Hence, if we impose the constraint (61), relating $J$ with $G$, and regard $H$ and L-L as functionals of two independent functions, e.g. $G$ and $F$, the minimizations of $H$ and L-L should result in the same field configurations, since the surface term does not contribute to the functional derivatives.
However, when $H$ and L-L are regarded as the functionals of three independent functions, F,GF,\ G and $J$, we should be careful, since the extremum conditions $\frac{\delta H}{\delta\Phi}=0$ and $\frac{\delta(-L)}{\delta\Phi}=0$ for $\Phi=F,G,J$ yield different results. To be more concrete, $\frac{\delta H}{\delta\Phi}=0$ does not yield the solutions to the equations of motion, as we have already pointed out.
Since we are interested in the solutions to the equations of motion, our task seems to solve the gradient flow equations with respect to the three independent functions, $F,G$ and $J$, where L-L takes place of $H$. By the way, only in solving the gradient flow equation, by some technical reason in the numerical computation, we adopt $H(\rho),I(\rho)$ instead of $F(\rho),J(\rho)$, defined by
@@
Unfortunately, by solving the gradient flow equations (69) to (71) numerically , we have found that the function $I$ shows divergent behavior as $s$ increases. The origin of this problem seems to come from the fact that in the L-L the term $E_{i}^{2}$ has a negative metric, which, when L-L is written in the quadratic form of the small deviations $\delta\Phi\ (\Phi=I,G,H)$ from the stationary point $\Phi_{st}$ (for which the r.h.s.s of (69) to (71) vanish) as $\Phi=\Phi_{st}+\delta\Phi$, probably leads to a negative eigenvalue of the matrix to describe the quadratic form and therefore the instability of the stationary point.
In order to overcome this problem, since the problematic term $E_{i}^{2}$ is mainly described by $I$, we think about the possibility to evade the equation (69) and replace it by imposing the analytic Gauss law constraint (61) , now written in terms of $I$ as
@@
for the purpose to express $I$ in terms of $G$ and substitute such obtained $I$ in the r.h.s. of the remaining gradient flow equation (70). Nice thing of this method is once $I$ is enforced to be related to $G$ by use of the constraint, L-L should not have the problem of negative eigenvalue mentioned above. This is because, under the Gauss law constraint, L-L differs from $H$ just by a constant (as is seen in (67)), while $H$ is clearly positive definite and does not possess the problem of negative eigenvalue.
@@
$+\left\{H^{\prime}-\sin\mu I^{\prime}-\frac{H-\sin\mu I}{\rho}+\cos^{2}\mu\frac{G(G-2)}{\rho}\right\}^{2}+\frac{2\rho^{2}}{g_{4}^{2}v_{4}^{4}\cos^{2}\mu}V_{4}(H/\rho)\Bigg{]},$